OpenStudy (anonymous):

solve for x:2|2x-2|+4=20

OpenStudy (mathstudent55):

First, subtract 4 from both sides. What do you get?

OpenStudy (anonymous):

what numbers do you subtract 4 from?

OpenStudy (solomonzelman):

Like this,\[\huge\color{red}{ 2 |2 x -2|+4=20 }\]\[\huge\color{red}{ 2 |2 x -2|+4\color{blue} { -4 } =20 }\color{blue} { -4 } \]\[\huge\color{red}{ 2 |2 x -2|=16 }\]\[\huge\color{red}{ SEE? }\]

OpenStudy (solomonzelman):

then divide both sides by 2.

OpenStudy (mathstudent55):